OpenStudy (anonymous):

Choose the correct simplification of the expression a^6b^3/a^4b^2

OpenStudy (anonymous):

\[\frac{a^6\cdot b^3}{a^4 \cdot b^2} = \frac{a^6}{a^4} \cdot \frac{b^3}{b^2} = a^6 \cdot a^{(-4)} \cdot b^3 \cdot b^{(-2)} = a^{6-4} \cdot b^{3-2}\] Try identifying the rules used here and then make sure to remember them, you'll proably need them very often in the future.

OpenStudy (anonymous):

I actually didn't do the very last step (execute the substractions). But I'm confident you'll be able to do it yourself